OpenStudy (anonymous):

Which property is illustrated by this example? 85 + (75 + 25) = (85 + 75) + 25

OpenStudy (uri):

Options? :3

OpenStudy (anonymous):

associative distributive communtative

OpenStudy (uri):

Disributive..?

OpenStudy (anonymous):

what about this Which property is illustrated by this example? 7 + (3 + 2) = 7 + (2 + 3)

OpenStudy (anonymous):

communitive

OpenStudy (uri):

Ask on question in one thread,Close this,Thanks.

OpenStudy (anonymous):

associative

OpenStudy (anonymous):

Which property is illustrated by this example? 5(8+6)=5*8+5*6

OpenStudy (uri):

@Sunshine1802 Don't help him/her here anymore.

OpenStudy (anonymous):

why thats what the websites for

OpenStudy (uri):

One question in one thread,close this. and ask in a new one.